Determine the polydispersity index for polystyrene having the following molecular weight distributions:

Determine the polydispersity index for polystyrene
Answer #1

The polydispersity index is the ratio of Molecular weight averages.

Polydispersity Index (PDI ) = \tiny \sum Mw / \tiny \sum Mn

PDI = ( 5000 + 7000 + 9000 + 11000 + 13000) / ( 2+15+17+12+4)

= 45000 / 50

= 900

Answer ; (PDI ) = 900